# Behavior Foundation Model: Towards Next-Generation Whole-Body Control System of Humanoid Robots

### v1

@@ 5.2.1 Ethical Issues
Ethical issues consistently accompany the development of diverse foundation models [185, 186, 187], which involve biased or unlicensed data, racial discrimination, and uncontrollable behaviors, etc. For BFMs, training on non-diverse motion datasets may encode demographic biases such as favoring movements natural to specific age groups or body types, which then propagate into robotic behaviors and create embodied forms of discrimination. Meanwhile, privacy risks escalate beyond data memorization to movement analytics, where rehabilitation or performance data could leak sensitive health information through generated motions. The physical instantiation of BFMs introduces unprecedented risks: unlike purely digital models, misaligned BFMs might reproduce unsafe or socially harmful behaviors (e.g., aggressive gestures or exclusionary motions) with real-world consequences. While techniques like differential privacy and federated learning offer partial solutions, they struggle with the continuous motion data’s temporal nature. In summary, BFMs demand novel governance frameworks that address both data provenance and normativity of real-time behavior.
@@ 5.2.2 Safety Mechanism
While most current BFMs focus primarily on proprioceptive inputs, integrating multimodal information (e.g., visual, linguistic, and auditory cues) has emerged as a promising direction for more generalizable and situationally aware control. However, multimodality introduces new vulnerabilities. Adversaries can exploit inconsistencies across modalities, as seen in the well-known CLIP case where an apple image was misclassified as an ”iPod” due to an overlaid label [191]. Such cross-modal confusion can be especially dangerous when a BFM is adapted for a unimodal task but retains sensitivity to irrelevant signals from other modalities. These challenges underscore the need to develop robust safety mechanisms for BFMs. Future work could prioritize adversarial robustness, cross-modal consistency checks, and disentangling modality-specific information to ensure predictable, trustworthy robot behavior in open-world environments.
@@ 6 Conclusion
In this paper, we present a systematic preview of the behavior foundation model, an emerging yet transformative paradigm for humanoid whole-body control systems. By pre-training on large-scale and diverse humanoid behavior data, BFMs learn a broad behavior coverage that enables few-shot or zero-shot adaptation to broad downstream tasks, eliminating the need for resource-intensive task-specific training. We establish a comprehensive taxonomy categorizing BFM approaches into supervised and unsupervised frameworks, while demonstrating their real-world applicability across healthcare, gaming, and industrial domains. Furthermore, we identify key research opportunities in high-level ML system integration, post-training optimization, and standardized evaluation mechanisms that could accelerate BFM development.

### v2

@@ 5.2.1 Ethical Issues
Ethical issues consistently accompany the development of diverse foundation models [199, 200, 201], which involve biased or unlicensed data, racial discrimination, and uncontrollable behaviors, etc. For BFMs, training on non-diverse motion datasets may encode demographic biases such as favoring movements natural to specific age groups or body types, which then propagate into robotic behaviors and create embodied forms of discrimination. Meanwhile, privacy risks escalate beyond data memorization to movement analytics, where rehabilitation or performance data could leak sensitive health information through generated motions. The physical instantiation of BFMs introduces unprecedented risks: unlike purely digital models, misaligned BFMs might reproduce unsafe or socially harmful behaviors (e.g., aggressive gestures, or exclusionary motions) with real-world consequences. While techniques like differential privacy and federated learning offer partial solutions, they struggle with the continuous motion data’s temporal nature. In summary, BFMs demand novel governance frameworks that address both data provenance and normativity of real-time behavior.
@@ 5.2.2 Safety Mechanism
While most current BFMs focus primarily on proprioceptive inputs, integrating multimodal information (e.g., visual, linguistic, and auditory cues) has emerged as a promising direction for more generalizable and situationally aware control. However, multimodality introduces new vulnerabilities. Adversaries can exploit inconsistencies across modalities, as seen in the well-known CLIP case where an apple image was misclassified as an ”iPod” due to an overlaid label [205]. Such cross-modal confusion can be especially dangerous when a BFM is adapted for an unimodal task but retains sensitivity to irrelevant signals from other modalities. These challenges underscore the need to develop robust safety mechanisms for BFMs. Future work could prioritize adversarial robustness, cross-modal consistency checks, and disentangling modality-specific information to ensure predictable, trustworthy robot behavior in open-world environments.
@@ 6 Conclusion
In this paper, we present a systematic preview of the behavior foundation model, an emerging and transformative paradigm of humanoid whole-body control systems. By pre-training on large-scale and diverse humanoid behavior data, BFMs learn broad behavior priors that enable few-shot or zero-shot adaptation to diverse downstream tasks, eliminating the need for resource-intensive and task-specific retraining. We establish a comprehensive taxonomy categorizing BFM approaches into pre-training pipelines and adaptation approaches while demonstrating their real-world applicability across industrial, gaming, and healthcare domains. Furthermore, we highlight future research opportunities in multimodal BFMs, high-level ML system integration, post-training optimization, and standardized evaluation mechanisms that could accelerate BFM development.